# Quasi-fuchsian flows and the coupled vortex equations

### v2

@@ 1. Introduction
In 1992, Ghys introduced a remarkable class of Anosov flows. Given a closed oriented surface $M$ of genus $\geq 2$, let $g_{1}$ and $g_{2}$ be two metrics of constant curvature $-1$, and let $[g_{1}]_{\mathrm{Teich}}$ and $[g_{2}]_{\mathrm{Teich}}$ be the corresponding points in Teichmüller space ${\mathcal{T}}(M)$. Théorème B in [Ghy92] establishes the existence of a smooth, orientable Anosov foliation $\phi_{[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}}$ on the unit tangent bundle $T_{1}M$ of positive half-lines tangent to the surface $M$, with the following properties: any flow parametrizing $\phi_{[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}}$ has a smooth weak stable foliation that is $C^{\infty}$-conjugate to the weak stable foliation of $g_{1}$, and a smooth weak unstable foliation that is $C^{\infty}$-conjugate to the weak unstable foliation of $g_{2}$. Moreover, a volume form is preserved if and only if $[g_{1}]_{\mathrm{Teich}}=[g_{2}]_{\mathrm{Teich}}$. Ghys called these flows quasi-Fuchsian flows, as his construction is analogous to that of quasi-Fuchsian groups obtained by coupling two Fuchsian groups [Mas88, IX.F].
@@
One can develop intuition for quasi-Fuchsian flows by first considering the case where $g_{1}$ and $g_{2}$ are $C^{3}$-close. In this case, when considering their respective geodesic flows on $T_{1}M$, the weak stable foliation of $g_{1}$ is transverse to the weak unstable foliation of $g_{2}$. Their intersection defines the Anosov foliation $\phi_{[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}}$, and hence [Ghy92, Théorème B] can be viewed as a statement about extending this class of flows for arbitrary pairs of points in Teichmüller space. Ghys’ construction is based on introducing a new differentiable structure on $T_{1}M$ such that the geodesic flow of a fixed background hyperbolic metric $h$ admits a transverse $(\text{PSL}(2,\mathbb{R})\times\text{PSL}(2,\mathbb{R}),\mathbb{RP}^{1}\times\mathbb{RP}^{1})$-structure, with a holonomy naturally constructed by coupling the holonomies of $g_{1}$ and $g_{2}$ (see [BG84, Section 1] and [CC00, 1.2.C] for background on transverse structures). Then, using the fact that homeomorphic closed 3-manifolds are diffeomorphic, $\phi_{[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}}$ is obtained by pulling back the geodesic flow of $h$ via this abstract (i.e., not explicit) diffeomorphism. While this construction is elegant and well-suited for classification, it complicates certain computations, especially if one is interested—as we are—in explicitly computing resonant 1-forms, winding cycles of SRB measures, and helicity, as in [CP25].
@@
By [MP19, Remark 5.3], for any holomorphic differential $A$ of degree $m$, there is a unique metric $g$ in the conformal class $[g]_{\mathrm{conf}}$ that solves (1.2), which we will refer to as the Blaschke metric. The unique hyperbolic metric in this class will be denoted by $\sigma$, so that $[g]_{\mathrm{conf}}=[\sigma]_{\mathrm{conf}}$, but unless $A=0$, $g$ and $\sigma$ do not coincide. Hence, this class of flows is parametrized by the data $([g]_{\mathrm{conf}},A)$, i.e. by a Riemann surface structure on $M$ and a holomorphic differential of degree $m$. The group $\text{Diff}_{0}(M)$ of diffeomorphisms isotopic to the identity acts naturally on this data and we shall denote the quotient by ${\mathcal{T}}\mathrm{D}_{m}(M)$, the Teichmüller space of holomorphic differentials of degree $m$. For $m=2$ we can think of ${\mathcal{T}}\mathrm{D}_{2}(M)$ as the tangent (or cotangent) bundle of ${\mathcal{T}}(M)$, and in this case we will write ${\mathcal{T}}\mathrm{QD}(M)$ for ${\mathcal{T}}\mathrm{D}_{2}(M)$.
@@
Given two hyperbolic metrics $\sigma$ and $\rho$ on $M$, it is well-known that there exists a unique harmonic diffeomorphism $f:(M,\sigma)\to(M,\rho)$ isotopic to the identity. Moreover, the Hopf quadratic differential $(f^{*}\rho)^{2,0}$ is holomorphic (here the $(2,0)$-part of $f^{*}\rho$ is taken with respect to the complex structure derived from $[\sigma]_{\mathrm{conf}}$), and thus we have a map
@@
where $\text{QD}([\sigma]_{\mathrm{conf}})$ denotes the vector space of holomorphic quadratic differentials on the Riemann surface determined by $[\sigma]_{\mathrm{conf}}$. This map is studied in [Sam78, Wol89], where it is shown that it is a homeomorphism, see [Wol89, Theorem 3.1]. (Note that this map only depends on the conformal class $[\sigma]_{\mathrm{conf}}$, which is being used as a base point.)
@@
where $[\sigma]_{\mathrm{conf}}$ is a conformal class on $M$ and $A\in\mathrm{QD}([g]_{\mathrm{conf}})$, descends to a bijection $\mathcal{G}:{\mathcal{T}}\mathrm{QD}(M)\to{\mathcal{T}}(M)\times{\mathcal{T}}(M)$ such that if we define $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}):=\mathcal{G}([\sigma]_{\mathrm{conf}},A)$ and we let $g\in[\sigma]_{\mathrm{conf}}$ be the Blaschke metric, then:
@@
    The flow of $F$ preserves a volume form if and only if $[g_{1}]_{\mathrm{Teich}}=[g_{2}]_{\mathrm{Teich}}$, which holds if and only if $A=0$.
@@
  • The sum of Dirichlet energies associated with the pair $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ produces a unique global minimum on ${\mathcal{T}}(M)$ and hence a point $[\sigma]_{\mathrm{Teich}}\in{\mathcal{T}}(M)$, which, as explained in [Sch93], should be thought of as the midpoint in ${\mathcal{T}}(M)$ between $[g_{1}]_{\mathrm{Teich}}$ and $[g_{2}]_{\mathrm{Teich}}$ since the pair of harmonic maps thus obtained have opposite Hopf differentials, $A$ and $-A$.
@@
$-2\text{\rm Re}\,A+\left(1+|A|^{2}_{g}\right)g\in[g_{2}]_{\mathrm{Teich}}.$
Often it is more practical to describe the points $[g_{1}]_{\mathrm{Teich}}$ and $[g_{2}]_{\mathrm{Teich}}$ using the marked length spectrum. In the course of proving the main theorem, we will also show that for a free homotopy class $\mathfrak{c}$ in $M$ we have
@@
for all closed orbits $\zeta$ of $F$. By the Livšic theorem, $V\lambda$ is a coboundary, and hence $F$ preserves a volume form (since $V\lambda$ is the divergence of $F$). By Theorem 1.1, Item d, this implies that $A=0$, and thus $[g_{1}]_{\mathrm{Teich}}=[g_{2}]_{\mathrm{Teich}}$.
@@ Acknowledgements
We are grateful to Martin Leguil and Andrey Gogolev for discussions related to their recent work [GLH24]. GPP was supported by NSF grant DMS-2347868.
@@ 2.1. Geometry of the unit circle bundle
It is straightforward to see that $\mho_{k}\to M$ has the structure of a smooth line bundle; a smooth section of this bundle may be seen as a smooth function on $SM$. In fact if $\mathcal{K}:=(T_{\mathbb{C}}^{*}M)^{1,0}$ is the associated canonical bundle and $\mathcal{K}^{-1}:=(T_{\mathbb{C}}^{*}M)^{0,1}$, then for any $k\in\mathbb{Z}$, $\mho_{k}$ may be identified with the tensor power $\mathcal{K}^{\otimes k}$ (here, if $k<0$ we set $\mathcal{K}^{\otimes k}=(\mathcal{K}^{-1})^{\otimes|k|}$). For any $k\in\mathbb{Z}_{\geq 0}$ and $x\in M$, there are natural maps
@@ 2.2. Coupled vortex equations
(2.3) ¯$\bar{\partial}A=0,\quad K_{g}=-1+(m-1)|A|_{g}^{2}.$
(See [DW15, Section 5] for the relation with the classical vortex equations in gauge theory.) It is known that when (2.3) holds, the flow of $F:=X+\lambda V$ for
@@ 2.3. Harmonic maps and the Hopf differential
When $M$ is two-dimensional (and orientable), it easy to see from definition that $e(f)\,d\mathrm{vol}_{\sigma}$ as well as the Dirichlet energy only depend on the conformal class $[\sigma]_{\mathrm{conf}}$, and hence the harmonic map equation only depends on the complex structure on $M$. The quadratic differential $A:=(f^{*}\rho)^{2,0}$ is called the Hopf differential of $f$. As an important fact, a direct computation in local (isothermal) coordinates shows that if $f$ is harmonic, then $A$ is holomorphic, see [Wol89, pages 451 and 452] (and references therein). (Conversely, if $A$ is holomorphic, then $f$ is harmonic assuming also that the Jacobian of $f$ is non-zero.) In this paper, we are exclusively concerned with the case when the target $N=M$ and $\rho$ is a hyperbolic metric. Then, given the data $[\sigma]_{\mathrm{conf}}$ and $\rho$ on $M$, there exists a unique harmonic diffeomorphism $f_{\sigma,\rho}:=f:(M,[\sigma]_{\mathrm{conf}})\to(M,\rho)$ such that $f$ is isotopic to the identity (see [Jos17, Theorem 9.7.2]).
@@
(2.9) $f_{\varphi^{*}\sigma,\phi^{*}\rho}=\phi^{-1}\circ f_{\sigma,\rho}\circ\varphi,\quad E_{\varphi^{*}\sigma,\phi^{*}\rho}(f_{\varphi^{*}\sigma,\phi^{*}\rho})=E_{\sigma,\rho}(f),$
@@
where $\sigma$ and $\rho$ are arbitrary representatives of $[\sigma]_{\mathrm{Teich}}$ and $[\rho]_{\mathrm{Teich}}$, respectively.
@@
The map $\Phi_{[\sigma]_{\mathrm{conf}}}$ is known to be a homeomorphism by the results in [Sam78, Wol89], see [Wol89, Theorem 3.1]. (This gives in particular that ${\mathcal{T}}(M)$ is homeomorphic to $\mathbb{R}^{6g(M)-6}$ where $g(M)$ is the genus of $M$, since $66g(M)-$ 6 is the real dimension of the space of holomorphic quadratic differentials.) We can think of $\Phi^{-1}_{[\sigma]_{\mathrm{conf}}}$ as a kind of exponential map for Teichmüller space; given $A\in\text{QD}([\sigma]_{\mathrm{conf}})$ an explicit hyperbolic metric representing $\Phi_{[\sigma]_{\mathrm{conf}}}^{-1}(A)\in{\mathcal{T}}(M)$ is given in Lemma 4.1 below in terms of $A$ and the Blaschke metric in the class $[\sigma]_{\mathrm{conf}}$.
@@ 3. The bijection $\mathcal{G}$
This map has some good properties as shown in [FT87, Theorem 2.4] and [SY79] (see also references therein): it is smooth, proper and its derivative at $[\sigma]_{\mathrm{Teich}}$ is given by
@@
[…]
Fix a pair $[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}\in\mathcal{T}(M)$ and define the function $F:\mathcal{T}(M)\to\mathbb{R}_{>0}$
@@
Note that $F$ is also proper. Using (3.1) we see that $[\sigma]_{\mathrm{Teich}}$ is a critical point of $F$ if and only if
@@
Now since $F$ is proper and bounded below it has an absolute minimum $[\sigma_{\text{min}}]_{\mathrm{Teich}}\in\mathcal{T}(M)$. If we let $A:=\Phi_{[\sigma_{\text{min}}]_{\mathrm{conf}}}([g_{1}]_{\mathrm{Teich}})$, then we see that
@@
as desired. Note that the argument above shows (by definition of $\mathcal{G}$) that $\mathcal{G}([\sigma]_{\mathrm{conf}},A)=([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ if and only if $[\sigma]_{\mathrm{Teich}}$ is a critical point of $F$ and $A=\Phi_{[\sigma]_{\mathrm{conf}}}([g_{1}]_{\mathrm{Teich}})=-\Phi_{[\sigma]_{\mathrm{conf}}}([g_{2}]_{\mathrm{Teich}})$.
To show that $\mathcal{G}$ is injective it suffices to invoke the proof of [Sch93, Proposition 2.12] which shows that the global minimum $[\sigma_{\text{min}}]_{\mathrm{Teich}}$ is in fact the unique critical point of $F$. ∎
@@
It is possible to recast the bijection $\mathcal{G}$ and Lemma 4.1 using AdS geometry; this originates in the work of Mess in [Mes07]. A pair $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ uniquely determines a maximal global hyperbolic AdS structure on $M\times\mathbb{R}$. It turns out that such an AdS manifold contains a unique Cauchy surface that is maximal, that is, it has zero mean curvature [BBZ07]. In addition its first fundamental form matches precisely the Blaschke metric $g$; one may see this by checking that the coupled vortex equations agree with the Gauss-Codazzi equations for the maximal surface. Moreover, the maximal surface comes equipped with the two “projection” maps that are the unique harmonic diffeomorphisms isotopic to the identity that we described earlier having opposite Hopf differentials. This approach also delivers Lemma 4.1 as explained in [BS20, Section 7]. The bijection $\mathcal{G}$ may also be constructed using AdS geometry, see [KS07, Theorem 3.8] and [BS20, Subsection 7.3].
@@
Since $\text{id}+{\mathbb{A}}_{x}$ is a linear isometry, ±$\pm 1$ cannot be eigenvalues of ${\mathbb{A}}_{x}$ (recall that ${\mathbb{A}}_{x}$ has zero trace). This implies right away that $K_{g}<0$ or equivalently $|A|_{g}<1$ (thus obtaining a proof of (2.5) for $m=2$). Indeed, if there was a point in $M$ where $|A|_{g}>1$, since $A$ has $04g(M)-4>$ 0 zeroes with multiplicity (here $g(M)$ is the genus of $M$), by continuity we can also find a point where $|A|_{g}=1$; this is however handled by the previous argument. Hence the eigenvalues of $\text{\rm id}+{\mathbb{A}}_{x}$ are positive thus showing that it preserves orientation. ∎
@@
The determinant of the preceding matrix is equal to $01-\frac{1}{4}(V\lambda)^{2}-\lambda^{2}=-K_{g}>$ 0 where we used (2.3), (2.5), and (2.6), so $a=b=0$ as claimed. ∎
@@
We are now ready to put the pieces together and complete the proof of our main result. By Proposition 3.1 the map $\mathcal{G}$ is a bijection. If $\mathcal{G}([\sigma]_{\mathrm{conf}},A)=([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ by definition we know that $A$ is the Hopf differential of the unique harmonic diffeomorphism $f:(M,[\sigma]_{\mathrm{conf}})\to(M,g_{1})$ isotopic to the identity (and $-A$ is the Hopf differential of the unique harmonic diffeomorphism from $(M,[\sigma]_{\mathrm{conf}})$ to $(M,g_{2})$). Obviously $h=f^{*}g_{1}$ is isometric to $g_{1}$ and its geodesic flow has weak stable foliation conjugate to that of $g_{1}$ via the lift of $f$. Next we claim that the principal bundle isomorphism $\mathcal{I}:SM_{g}\to SM_{h}$ conjugates the weak foliation of $F=X+\lambda V$ with the weak foliation of $X_{1}$, the geodesic vector field of $h$. Indeed Lemma 4.5 implies

### v3

@@ 1. Introduction
In 1992, Ghys introduced a remarkable class of Anosov flows. Given a closed oriented surface $M$ of genus $\geq 2$, let $g_{1}$ and $g_{2}$ be two metrics of constant curvature 1-1, and let $g_{1}]_{\mathrm{Teich}}$ and $g_{2}]_{\mathrm{Teich}}$ be the corresponding points in Teichmüller space ${\mathcal{T}}(M)$. Théorème B in [Ghy92] establishes the existence of a smooth, orientable Anosov foliation $\phi_{[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}}$ on the unit tangent bundle $T_{1}M$ of positive half-lines tangent to the surface $M$, with the following properties: any flow parametrizing $\phi_{[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}}$ has a smooth weak stable foliation that is $C^{\infty}$-conjugate to the weak stable foliation of $g_{1}$, and a smooth weak unstable foliation that is $C^{\infty}$-conjugate to the weak unstable foliation of $g_{2}$. Moreover, a volume form is preserved if and only if $g_{1}]_{\mathrm{Teich}}=[g_{2}]_{\mathrm{Teich}}$. Ghys called these flows quasi-Fuchsian flows, as his construction is analogous to that of quasi-Fuchsian groups obtained by coupling two Fuchsian groups [Mas88, IX.F].
@@
One can develop intuition for quasi-Fuchsian flows by first considering the case where $g_{1}$ and $g_{2}$ are $C^{3}$-close. In this case, when considering their respective geodesic flows on $T_{1}M$, the weak stable foliation of $g_{1}$ is transverse to the weak unstable foliation of $g_{2}$. Their intersection defines the Anosov foliation $\phi_{[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}}$, and hence [Ghy92, Théorème B] can be viewed as a statement about extending this class of flows for arbitrary pairs of points in Teichmüller space. Ghys’ construction is based on introducing a new differentiable structure on $T_{1}M$ such that the geodesic flow of a fixed background hyperbolic metric $h$ admits a transverse $(\text{PSL}(2,\mathbb{R})\times\text{PSL}(2,\mathbb{R}),\mathbb{RP}^{1}\times\mathbb{RP}^{1})$-structure, with a holonomy naturally constructed by coupling the holonomies of $g_{1}$ and $g_{2}$ (see [BG84, Section 1] and [CC00, 1.2.C] for background on transverse structures). Then, using the fact that homeomorphic closed 3-manifolds are diffeomorphic, $\phi_{[g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}}$ is obtained by pulling back the geodesic flow of $h$ via this abstract (i.e., not explicit) diffeomorphism. While this construction is elegant and well-suited for classification, it complicates certain computations, especially if one is interested—as we are—in explicitly computing resonant 1-forms, winding cycles of SRB measures, and helicity, as in [CP25] (for background on the microlocal approach to hyperbolic flows see [Lef25]).
@@
By [MP19, Remark 5.3], for any holomorphic differential $A$ of degree $m$, there is a unique metric $g$ in the conformal class $g]_{\mathrm{conf}}$ that solves (1.2), which we will refer to as the Blaschke metric. The unique hyperbolic metric in this class will be denoted by $\sigma$, so that $g]_{\mathrm{conf}}=[\sigma]_{\mathrm{conf}}$, but unless $A=0$, $g$ and $\sigma$ do not coincide. Hence, this class of flows is parametrized by the data $([g]_{\mathrm{conf}},A)$, i.e. by a Riemann surface structure on $M$ and a holomorphic differential of degree $m$. The group $\text{Diff}_{0}(M)$ of diffeomorphisms isotopic to the identity acts naturally on this data and we shall denote the quotient by ${\mathcal{T}}\mathrm{D}_{m}(M)$, the Teichmüller space of holomorphic differentials of degree $m$. For $m=2$ we can think of ${\mathcal{T}}\mathrm{D}_{2}(M)$ as the tangent (or cotangent) bundle of ${\mathcal{T}}(M)$, and in this case we will write ${\mathcal{T}}\mathrm{QD}(M)$ for ${\mathcal{T}}\mathrm{D}_{2}(M)$.
@@
Given two hyperbolic metrics $\sigma$ and $\rho$ on $M$, it is well-known that there exists a unique harmonic diffeomorphism $f:(M,\sigma)\to(M,\rho)$ isotopic to the identity. Moreover, the Hopf quadratic differential $(f^{*}\rho)^{2,0}$ is holomorphic (here the $(2,0)$-part of $f^{*}\rho$ is taken with respect to the complex structure derived from $\sigma]_{\mathrm{conf}}$), and thus we have a map
@@
where $\text{QD}([\sigma]_{\mathrm{conf}})$ denotes the vector space of holomorphic quadratic differentials on the Riemann surface determined by $\sigma]_{\mathrm{conf}}$. This map is studied in [Sam78, Wol89], where it is shown that it is a homeomorphism, see [Wol89, Theorem 3.1]. (Note that this map only depends on the conformal class $\sigma]_{\mathrm{conf}}$, which is being used as a base point.)
@@
where $\sigma]_{\mathrm{conf}}$ is a conformal class on $M$ and $A\in\mathrm{QD}([\sigma]_{\mathrm{conf}})$, descends to a bijection $\mathcal{G}:{\mathcal{T}}\mathrm{QD}(M)\to{\mathcal{T}}(M)\times{\mathcal{T}}(M)$ such that if we define $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}):=\mathcal{G}([\sigma]_{\mathrm{conf}},A)$ and we let $g\in[\sigma]_{\mathrm{conf}}$ be the Blaschke metric, then:
@@
    The flow of $F$ preserves a volume form if and only if $g_{1}]_{\mathrm{Teich}}=[g_{2}]_{\mathrm{Teich}}$, which holds if and only if $A=0$.
@@
  • The sum of Dirichlet energies associated with the pair $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ produces a unique global minimum on ${\mathcal{T}}(M)$ and hence a point $\sigma]_{\mathrm{Teich}}\in{\mathcal{T}}(M)$, which, as explained in [Sch93], should be thought of as the midpoint in ${\mathcal{T}}(M)$ between $g_{1}]_{\mathrm{Teich}}$ and $g_{2}]_{\mathrm{Teich}}$ since the pair of harmonic maps thus obtained have opposite Hopf differentials, $A$ and A-A.
@@
$2\text{\rm Re}\,A+\left(1+|A|^{2}_{g}\right)g\in[g_{2}]_{\mathrm{Teich}}.$
Often it is more practical to describe the points $g_{1}]_{\mathrm{Teich}}$ and $g_{2}]_{\mathrm{Teich}}$ using the marked length spectrum. In the course of proving the main theorem, we will also show that for a free homotopy class $\mathfrak{c}$ in $M$ we have
@@
for all closed orbits $\zeta$ of $F$. By the Livšic theorem, $V\lambda$ is a coboundary, and hence $F$ preserves a volume form (since $V\lambda$ is the divergence of $F$). By Theorem 1.1, Item d, this implies that $A=0$, and thus $g_{1}]_{\mathrm{Teich}}=[g_{2}]_{\mathrm{Teich}}$.
In addition to the applications in [CP25] mentioned above, we may raise also a few interesting questions related to our approach. Namely, quasi-Fuchsian flows associate an Anosov flow on a $3$-manifold to a pair of co-compact representations $\pi_{1}(M)\to\mathrm{PSL}(2,\mathbb{R})$; in a parallel development, from the same data [DGM25] (see also [GBLW26]) construct an Axiom A flow from the spacelike geodesic flow on the $3$-dimensional Anti-de Sitter space, with the same periods of closed orbits. It is natural to ask what is the precise relation between these two flows. Another question is to clarify semisimplicity for resonant forms and more generally the behaviour of the Ruelle zeta function for quasi-Fuchsian flows [CP25]; we do this for cubic differentials in [CP26].
@@ Acknowledgements
We are grateful to Martin Leguil and Andrey Gogolev for discussions related to their recent work [GLH24]. GPP was supported by NSF grant DMS-2347868. MC received funding from an Ambizione grant (project number 201806) from the Swiss National Science Foundation.
@@ 2.1. Geometry of the unit circle bundle
It is straightforward to see that $\mho_{k}\to M$ has the structure of a smooth line bundle; a smooth section of this bundle may be seen as a smooth function on $SM$. In fact if $\mathcal{K}:=(T_{\mathbb{C}}^{*}M)^{1,0}$ is the associated canonical bundle and $\mathcal{K}^{-1}:=(T_{\mathbb{C}}^{*}M)^{0,1}$, then for any $\in\mathbb{Z}, k\mho_{k}$ may be identified with the tensor power $\mathcal{K}^{\otimes k}$ (here, if $k<0$ we set $\mathcal{K}^{\otimes k}=(\mathcal{K}^{-1})^{\otimes|k|}$). For any $k\in\mathbb{Z}_{\geq 0}$ and $x\in M$, there are natural maps
@@ 2.2. Coupled vortex equations
(2.3) $\bar{\partial}A=0,\quad K_{g}=-1+(m-1)|A|_{g}^{2}.$
(See [Li19] for the relation with the classical vortex equations in gauge theory.) It is known that when (2.3) holds, the flow of $F:=X+\lambda V$ for
@@ 2.3. Harmonic maps and the Hopf differential
When $M$ is two-dimensional (and orientable), it is easy to see from the definition that $e(f)\,d\mathrm{vol}_{\sigma}$ as well as the Dirichlet energy only depend on the conformal class $\sigma]_{\mathrm{conf}}$, and hence the harmonic map equation only depends on the complex structure on $M$. The quadratic differential $A:=(f^{*}\rho)^{2,0}$ is called the Hopf differential of $f$. As an important fact, a direct computation in local (isothermal) coordinates shows that if $f$ is harmonic, then $A$ is holomorphic, see [Wol89, pages 451 and 452] (and references therein). (Conversely, if $A$ is holomorphic, then $f$ is harmonic assuming also that the Jacobian of $f$ is non-zero.) In this paper, we are exclusively concerned with the case when the target $N=M$ and $\rho$ is a hyperbolic metric. Then, given the data $\sigma]_{\mathrm{conf}}$ and $\rho$ on $M$, there exists a unique harmonic diffeomorphism $f_{\sigma,\rho}:=f:(M,[\sigma]_{\mathrm{conf}})\to(M,\rho)$ such that $f$ is isotopic to the identity (see [Jos17, Theorem 9.7.2]).
@@
(2.9) $f_{\varphi^{*}\sigma,\phi^{*}\rho}=\phi^{-1}\circ f_{\sigma,\rho}\circ\varphi,\quad E_{\varphi^{*}\sigma,\phi^{*}\rho}(f_{\varphi^{*}\sigma,\phi^{*}\rho})=E_{\sigma,\rho}(f).$
@@
where $\sigma$ and $\rho$ are arbitrary representatives of $\sigma]_{\mathrm{Teich}}$ and $\rho]_{\mathrm{Teich}}$, respectively.
@@
The map $\Phi_{[\sigma]_{\mathrm{conf}}}$ is known to be a homeomorphism by the results in [Sam78, Wol89], see [Wol89, Theorem 3.1]. (This gives in particular that ${\mathcal{T}}(M)$ is homeomorphic to $\mathbb{R}^{6g(M)-6}$ where $g(M)$ is the genus of $M$, since 6g(M)66g(M)-6 is the real dimension of the space of holomorphic quadratic differentials.) We can think of $\Phi^{-1}_{[\sigma]_{\mathrm{conf}}}$ as a kind of exponential map for Teichmüller space; given $A\in\text{QD}([\sigma]_{\mathrm{conf}})$ an explicit hyperbolic metric representing $\Phi_{[\sigma]_{\mathrm{conf}}}^{-1}(A)\in{\mathcal{T}}(M)$ is given in Lemma 4.1 below in terms of $A$ and the Blaschke metric in the class $\sigma]_{\mathrm{conf}}$.
@@ 3. The bijection $\mathcal{G}$
This map has some good properties as shown in [FT87, Theorem 2.4] and [SY79] (see also references therein): it is smooth, proper and its derivative at $\sigma]_{\mathrm{Teich}}$ is given by
@@
where $\eta\in T_{[\sigma]_{\mathrm{Teich}}}\mathcal{T}(M)$ (identified with $\mathrm{QD}([\sigma]_{\mathrm{conf}})$), and $\xi=\mathrm{Re}(\varphi)$ with $\varphi=\Phi_{[\sigma]_{\mathrm{conf}}}([g_{0}]_{\mathrm{Teich}})$ is the Hopf differential of the unique harmonic map defined by $\sigma$ and $g_{0}$. The inner product $\langle{\bullet,\bullet}\rangle$ is given by the Weil-Petersson metric ($L^{2}$).
Fix a pair $g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}\in\mathcal{T}(M)$ and define the function $F:\mathcal{T}(M)\to\mathbb{R}_{>0}$
@@
Note that $F$ is also proper. Using (3.1) we see that $\sigma]_{\mathrm{Teich}}$ is a critical point of $F$ if and only if
@@
Now since $F$ is proper and bounded below it has an absolute minimum $\sigma_{\text{min}}]_{\mathrm{Teich}}\in\mathcal{T}(M)$. If we let $A:=\Phi_{[\sigma_{\text{min}}]_{\mathrm{conf}}}([g_{1}]_{\mathrm{Teich}})$, then we see that
@@
as desired. Note that the argument above shows (by definition of $\mathcal{G}$) that $\mathcal{G}([\sigma]_{\mathrm{conf}},A)=([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ if and only if $\sigma]_{\mathrm{Teich}}$ is a critical point of $F$ and $A=\Phi_{[\sigma]_{\mathrm{conf}}}([g_{1}]_{\mathrm{Teich}})=-\Phi_{[\sigma]_{\mathrm{conf}}}([g_{2}]_{\mathrm{Teich}})$.
To show that $\mathcal{G}$ is injective it suffices to invoke the proof of [Sch93, Proposition 2.12] which shows that the global minimum $\sigma_{\text{min}}]_{\mathrm{Teich}}$ is in fact the unique critical point of $F$. ∎
@@
It is possible to recast the bijection $\mathcal{G}$ and Lemma 4.1 using AdS geometry; this originates in the work of Mess in [Mes07]. A pair $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ uniquely determines a maximal global hyperbolic AdS structure on $M\times\mathbb{R}$. It turns out that such an AdS manifold contains a unique Cauchy surface that is maximal, that is, it has zero mean curvature [BBZ07]. In addition its first fundamental form matches precisely the Blaschke metric $g$; one may see this by checking that the coupled vortex equations agree with the Gauss-Codazzi equations for the maximal surface. Moreover, the maximal surface comes equipped with the two “projection” maps that are the unique harmonic diffeomorphisms isotopic to the identity that we described earlier having opposite Hopf differentials. This approach also delivers Lemma 4.1 as explained in [BS20, Section 7]. The bijection $\mathcal{G}$ may also be constructed using AdS geometry, see [KS07, Theorem 3.8] and [BS10].
@@
Since $\text{id}+{\mathbb{A}}_{x}$ is a linear isometry, $\pm 1$ cannot be eigenvalues of ${\mathbb{A}}_{x}$ (recall that ${\mathbb{A}}_{x}$ has zero trace). This implies right away that $K_{g}<0$ or equivalently $|A|_{g}<1$ (thus obtaining a proof of (2.5) for $m=2$). Indeed, if there was a point in $M$ where $|A|_{g}>1$, since $A$ has 4g(M)4>04g(M)-4>0 zeroes with multiplicity (here $g(M)$ is the genus of $M$), by continuity we can also find a point where $|A|_{g}=1$; this is however handled by the previous argument. Hence the eigenvalues of $\text{\rm id}+{\mathbb{A}}_{x}$ are positive thus showing that it preserves orientation. ∎
@@
The determinant of the preceding matrix is equal to $1-\frac{1}{4}(V\lambda)^{2}-\lambda^{2}=-K_{g}>0$ where we used (2.3), (2.5), and (2.6), so $a=b=0$ as claimed. ∎
@@
We are now ready to put the pieces together and complete the proof of our main result. By Proposition 3.1 the map $\mathcal{G}$ is a bijection. If $\mathcal{G}([\sigma]_{\mathrm{conf}},A)=([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ by definition we know that $A$ is the Hopf differential of the unique harmonic diffeomorphism $f:(M,[\sigma]_{\mathrm{conf}})\to(M,g_{1})$ isotopic to the identity (and A-A is the Hopf differential of the unique harmonic diffeomorphism from $(M,[\sigma]_{\mathrm{conf}})$ to $(M,g_{2})$). Obviously $h=f^{*}g_{1}$ is isometric to $g_{1}$ and its geodesic flow has weak stable foliation conjugate to that of $g_{1}$ via the lift of $f$. Next we claim that the principal bundle isomorphism $\mathcal{I}:SM_{g}\to SM_{h}$ conjugates the weak foliation of $F=X+\lambda V$ with the weak foliation of $X_{1}$, the geodesic vector field of $h$. Indeed Lemma 4.5 implies